# Improving Human Verification of LLM Reasoning through Interactive Explanation Interfaces

### v1

@@ Abstract
The reasoning capabilities of Large Language Models (LLMs) have led to their increasing employment in several critical applications, particularly education, where they support problem-solving, tutoring, and personalized study. While there are a plethora of works showing the effectiveness of LLMs in generating step-by-step solutions through chain-of-thought (CoT) reasoning on reasoning benchmarks, little is understood about whether the generated CoT is helpful for end-users in improving their ability to comprehend mathematical reasoning problems and detect errors/hallucinations in LLM-generated solutions. To address this gap and contribute to understanding how reasoning can improve human-AI interaction, we present three new interactive reasoning interfaces: interactive CoT (iCoT), interactive Program-of-Thought (iPoT), and interactive Graph (iGraph), and a novel framework that generates the LLM’s reasoning from traditional CoT to alternative, interactive formats. Across 125 participants, we found that interactive interfaces significantly improved performance. Specifically, the iGraph interface yielded the highest clarity and error detection rate ($85.6\%$), followed by iPoT ($82.5\%$), iCoT ($80.6\%$), all outperforming standard CoT ($73.5\%$). Interactive interfaces also led to faster response times, where participants using iGraph were fastest ($57.9$ secs), compared to iCoT and iPoT ($60$ secs), and the standard CoT baseline ($64.7$ secs). Furthermore, participants preferred the iGraph reasoning interface, citing its superior ability to enable users to follow the LLM’s reasoning process. We discuss the implications of these results and provide recommendations for the future design of reasoning models. Experiment interfaces are available on Huggingface.
@@ 1 Introduction
Large language models (LLMs) are now woven into high-stakes, real-world workflows, including education, where they assist with problem solving, tutoring, and personalized study [1, 2, 3]. A key driver of their recent success on complex tasks is chain-of-thought (CoT) prompting [4]: eliciting step-by-step intermediate reasoning often yields dramatic accuracy gains on arithmetic, commonsense, and symbolic reasoning benchmarks, with state-of-the-art performance on GSM8K and related math datasets. However, despite these advancements, the current format for presenting LLM explanations is far from optimal [5].
@@
Despite these advances, it remains unclear whether text-only CoT explanations actually help end users—students, teachers, or non-expert verifiers—understand and check mathematical reasoning. While these outputs contain rich information, their format introduces significant challenges for human interpretation [7, 8, 9]. In practice, CoT often appears as a long “wall of text” that can be time-consuming, obscure structure, increase cognitive load, and make error detection harder precisely when it matters. For example, when an explanation contains a subtle mistake in one intermediate step, detecting that error often requires carefully parsing multiple sentences and mentally reconstructing the reasoning flow [10]. More broadly, the explainable AI (XAI) literature shows that explanation techniques are frequently designed around what researchers can produce rather than what end users need; users instead report wanting practically useful information that improves collaboration, trust calibration, and feedback loops with AI systems[11, 12, 13]. This gap motivates rethinking how reasoning is surfaced to end-users in contexts like education, where clarity, verification, and engagement are essential.
@@
RQ1) Do interactive explanation formats improve verification accuracy compared to traditional text outputs?
RQ2) How do these formats influence the time and effort required to verify reasoning?

### v2

@@ Abstract
The reasoning capabilities of Large Language Models (LLMs) have led to their increasing employment in several critical applications, particularly education, where they support problem-solving, tutoring, and personalized study. Chain-of-thought (CoT) reasoning capabilities [1, 2] are well-known to help LLMs decompose a problem into steps and explore the solution spaces more effectively, leading to impressive performance on mathematical and reasoning benchmarks. As the length of CoT tokens per question increases substantially to even thousands of tokens per question [1], it is unknown how users could comprehend LLM reasoning and detect errors or hallucinations. To address this problem and understand how reasoning can improve human-AI interaction, we present three new interactive reasoning interfaces: interactive CoT (iCoT), interactive Program-of-Thought (iPoT), and interactive Graph (iGraph). That is, we ask LLMs themselves to generate an interactive web interface wrapped around the original CoT content, which may be presented in text (iCoT), graphs (iGraph) or code (iPoT). This interface allows users to interact with and provide a novel experience in reading and validating the reasoning chains of LLMs. Across a study of 125 participants, interactive interfaces significantly improve user performance. Specifically, iGraph users score the highest error detection rate ($85.6\%$), followed by iPoT ($82.5\%$), iCoT ($80.6\%$), all outperforming standard CoT ($73.5\%$). Interactive interfaces also lead to faster user validation time—iGraph users are faster ($57.9$ secs per question) than the users of iCoT and iPoT ($60$ secs) and the standard CoT ($64.7$ secs). A post-study questionnaire shows that users prefer iGraph, citing its superior ability to enable them to follow the LLM’s reasoning. We discuss the implications of these results and provide recommendations for the future design of reasoning models. Try our interactive interfaces on HuggingFace.
@@ 1 Introduction
Large language models (LLMs) are now woven into high-stakes, real-world workflows, including education, where they assist with problem solving, tutoring, and personalized study [3, 4, 5]. A key driver of their recent success on complex tasks is chain-of-thought (CoT) prompting [6]: eliciting step-by-step intermediate reasoning often yields dramatic accuracy gains on arithmetic, commonsense, and symbolic reasoning benchmarks, with state-of-the-art performance on GSM8K and related math datasets. However, the current format for presenting LLM explanations may be far from optimal [7].
@@
While CoT outputs contain rich information, their long-text presentation introduces significant challenges for human interpretation [9, 10, 11]. In practice, especially in modern reasoning LLMs [1], CoT often appears as a long “wall of text” that can be time-consuming, obscure structure, increase cognitive load, and make error detection harder precisely when it matters. For example, when an explanation contains a subtle mistake in one intermediate step, detecting that error often requires carefully parsing multiple sentences and mentally reconstructing the reasoning flow [12]. More broadly, the explainable AI (XAI) literature shows that explanation techniques are frequently designed around what researchers can produce rather than what end users need; users instead report wanting practically useful information that improves collaboration, trust calibration, and feedback loops with AI systems [13, 14, 15]. This gap motivates rethinking how reasoning is surfaced to end-users in contexts like education, where clarity, verification, and engagement are essential.
@@
RQ1) Do interactive explanation formats help users improve verification accuracy compared to traditional text outputs?
RQ2) How do these formats influence the time and effort required for users to verify reasoning?